# Symmetric square type L–series

### v1

@@ 3. Modular forms
Let $k\in(\mathbb{Z}/2)\setminus\mathbb{Z}$ with $k\geq 5/2$, $N\in\mathbb{N}$, and $\chi$ be a character modulo $4N$. Let
@@
be in $\mathcal{S}_{k}(\Gamma_{0}(4N),\chi)$. Then for $t\in\mathbb{N}$ and $n\nmid(2N)^{\infty}$ we find
@@
The second bound goes back to Iwaniec [Iwaniew1987] and Duke [Duke1988] in a special case and has been generalised by Waibel [Waibel2018, Thm. 1 p. 186].
@@ 3.1. Eisenstein series
Beyond the theta series presented above, Eisenstein series represent another important class of modular forms. Our analysis requires the use of non-holomorphic Eisenstein series, which we introduce briefly based upon [Kiefer2021, 2.5 pp. 30-36]. The statements are essential for our investigations, more specifically for deriving meromorphic continuation of the symmetric square type $L$-functions arising from the Rankin–Selberg type integrals in Section 5. In the following, let $(L,\operatorname{q})$ be a non-degenerate quadratic lattice and $k\in\tfrac{1}{2}\mathbb{Z}$ be a half integer.
@@
The Eisenstein series $E_{L,\lambda,k}$ converges normally on $\mathbb{H}$ for $\Re(s)>1-\frac{k}{2}$, is real analytic in $\tau$ and an eigenfunction of the hyperbolic Laplace operator of weight $k$ with Eigenvalue $s(s+k-1)$.
@@
$\mathcal{W}_{s}(v)\coloneq\lvert v\rvert^{-k/2}\mathcal{W}_{\operatorname{sgn}(v)k/2,(1-k)/2-s}(\lvert v\rvert),$
where $\mathcal{W}_{\kappa,\mu}(z)$ is the usual Whittaker function as in [OlverClark2010, (13.14.3) p. 334].
Proposition 3.16 ([BruinierKuehn2003, Prop. 3.1 p. 1695]).
@@
$\begin{cases}(\delta_{\lambda,\mu}+i^{\operatorname{sig}(\mathcal{L})+2k}\delta_{-\lambda,\mu})v^{s}+2\pi v^{1-k-s}\frac{\Gamma(k+2s-1)}{\Gamma(k+s)\Gamma(s)}\cdot\sum_{x\in\mathbb{Z}\setminus\{0\}}\lvert 2c\rvert^{1-k-2s}H_{c}(\beta,0,\mu,0),&\textnormal{ if }n=0,\\ \frac{2^{k}\pi^{s+k-1}}{\Gamma(s+k)}\mathcal{W}_{s}(4\pi nv)\cdot\sum_{c\in\mathbb{Z}\setminus\{0\}}\lvert c\rvert^{1-k-2s}H_{c}(\lambda,0,\mu,n),&\textnormal{ if }n>0,\\ \frac{2^{k}\pi^{s+k-1}}{\Gamma(s)}\mathcal{W}_{s}(4\pi nv)\cdot\sum_{c\in\mathbb{Z}\setminus\{0\}}\lvert c\rvert^{1-k-2s}H_{c}(\lambda,0,\mu,n),&\textnormal{ if }n<0.\\ \end{cases}$
@@
According to [Hejhal1983, p. 372] for an isotropic element $\lambda\in\mathcal{L}$
@@
From the Fourier expansion of $E_{L,\lambda,k}$ and the properties of the Whittaker function $\mathcal{W}_{s}$ (cf. [OlverClark2010, p. 335 (13.14.21)]) follows an asymptotic bound for the Eisenstein series which is relevant for the application of the Rankin-Selberg method.
@@
According to [Hejhal1983, p. 372] we have the following transformation property.
@@ 4. Hecke theory
Note that this result is consistent with [Miyake2006, Lem. 4.5.7 (1) p. 140], where $T(p,p)=T_{p\mathcal{I}}$ and $T(1,p^{e})=\mathcal{T}(p^{e})$ in our notation. Note also that the number $N$ in Miyake may be assumed to equal $1$ in the current setting. By utilising the statement above, we may derive a relation for square indices which is the relevant case in the vector-valued setting.
@@
    =\;= ${\mathcal{T}(p)}\left[\mathcal{T}(p)\mathcal{T}(p^{r})-p\cdot T_{p\mathcal{I}}\mathcal{T}(p^{r-1})\right]-p\cdot T_{p\mathcal{I}}\mathcal{T}(p^{r})$
    =\;= $\left[\mathcal{T}(p^{2})+(p+1)\cdot T_{p\mathcal{I}}\right]\mathcal{T}(p^{r})$
@@
    =\;= $\left[\mathcal{T}(p^{2})+(1-p)\cdot T_{p\mathcal{I}}\right]\mathcal{T}(p^{r})-(1+\delta_{r=2}p^{-1})\left(p\cdot T_{p\mathcal{I}}\right)^{2}\mathcal{T}(p^{r-2}).$
The Hecke operators introduced above not only constitute an abstract algebra but also operate on spaces of modular forms. Their action on Fourier coefficients of modular forms facilitates the derivation of product expansions for associated $L$-series, which motivates their examination in this work. The theory for scalar-valued modular forms is well known, however, naively extending the explicit construction of Hecke operators for the scalar-valued case to vector-valued modular forms fails, due to the Weil representation ${\rho}_{L}$ not being extendable to $\operatorname{GL}_{2}(\mathbb{Z}/\operatorname{lev}(L)\mathbb{Z})$ as a linear action. An alternative construction circumventing this issue is presented in [BruinierStein] and is consistent with classical Hecke operators in the scalar-valued case. We will briefly sketch the content necessary for our proceedings.
@@
That the notation above is well-defined is non-trivial (cf.  [BruinierStein, Sec. 5]).
@@
    In the general case, by the contravariance of the slash operator (cf. Definition 4.3), and invariance of $f$ and $h$ under $\overline{\Gamma(1)}$, we may reduce to the choice $\alpha=g(m^{2})$. This, however is solved by the same argument as above combined with Lemma 4.5 in the components of the group algebra. ∎
@@
For any $M\in\mathbb{N}$, the algebra $\mathcal{H}_{M}^{2}$ acts on $\mathcal{S}_{L,k}$ as a commutative self-adjoint algebra of cusp form preserving linear operators.
@@
Now that the existence of eigenforms is certain, a bound on the respective eigenvalues was advantageous. An effective method for obtaining such bounds was discovered by Kohnen [Kohnen1987] even though the bound itself, at least for scalar-valued Siegel Modular forms, had been known before (cf. [Weissauer1984, Sec. 6]).
@@
    =\;= $\sum_{\lambda,\mu\in\mathcal{L}}\langle f_{\lambda}|_{k}g(m^{2})g^{\ast}(m^{2}),f_{\mu}|_{k}g(m^{2})g^{\ast}(m^{2})\rangle\langle\mathfrak{e}_{\lambda}|_{k}g(m^{2})g^{\ast}(m^{2}),\mathfrak{e}_{\mu}|_{k}g(m^{2})g^{\ast}(m^{2})\rangle$
    =\;= $\sum_{\lambda,\mu\in\mathcal{L}}\langle f_{\lambda},f_{\mu}\rangle\lvert{{}_{m}\mathcal{L}}\rvert^{2}\langle\mathfrak{e}_{\lambda},\mathfrak{e}_{\mu}\rangle$
    =\;= $\lvert{{}_{m}\mathcal{L}}\rvert^{2}\|f\|_{2}^{2}.$
@@
    Finally, for the choice $m=p$, there are $p(p+1)$ cosets in the class $\overline{\Gamma(1)}\backslash\overline{\Gamma(1)}g(m^{2})\overline{\Gamma(1)}$ (cf. [ShimuraMF_HalfIntegralWeight, p. 451]), finishing the proof. ∎
@@ 4.1. Hecke action on Fourier coefficients
Note that for our purposes a formula that is valid for all primes is required. An advance in that direction has been made in [SteinHecke, Thm. 5.4 p. 246] for all odd primes. However, it has been pointed out by [Bouchard2018] that the cited Theorem is flawed. The original author has submitted a correction [Stein2021], representing the previously arising, but flawed, coefficients in a more abstract fashion in terms of certain representation numbers of the lattice that also appear as coefficients of Eisenstein series in [BruinierKuss2001, p. 447].
These representation numbers, however, are not explicit enough to derive product expansions of $L$-series. Hence, the coefficients have to be calculated more explicitly based on the exposition [SteinHecke]. To this end we have to let $\mathcal{T}(p^{2})$ explicitly act on $f\in\mathcal{M}_{L,k}$ where the operation is given by letting right cosets of
@@
operate as in Definition 4.4. Recall that by [BruinierStein, p. 263] or [ShimuraMF_HalfIntegralWeight, p. 451] we find
@@
The operation on the scalar component functions of the vector-valued modular form via the Petersson slash operator is well known and understood. As a consequence, it suffices to describe the action of the elements in (4.18) on the discriminant group. The action of $\gamma_{b}$ is presented in [SteinHecke, Thm. 5.1 p. 243] in greater generality than required.
@@
The action of $\beta_{h}$, however, will be extracted from [Bouchard2018, Prop. 5.3 p. 26] where we select the special case of $l=a=1$.
@@
where $G_{L,p}(n,h)=\sum_{v\in L/p^{n}L}e\left(\frac{h}{p^{n}}\operatorname{q}(v)\right)$ is the Gauss sum defined in Definition 2.4, $\ell_{\lambda/p}\in L^{\prime}$ is some element, such that $p\cdot\overline{\ell_{\lambda/p}}=\lambda$, ${{}^{p}\!\mathcal{L}}$ are the $p$-multiples in $\mathcal{L}$ and $\delta$ denotes a Kronecker symbol.
  • Proof:
@@
    where $\operatorname{q}(\mu)$ is to be understood as the quadratic form on $L^{\prime}$ of some preimage in $L^{\prime}$ of $\mu\in L^{\prime}/pL$. In order for the sum above not to vanish, it is necessary that $\lambda\in{{}^{p}\!\mathcal{L}}$ (cf. [Barnard2003, Lemma 5.2.1]). Note that by the isomorphism $(L^{\prime}/pL)/(L/pL)\simeq\mathcal{L}$, the solutions in $\mu$ of $\mu=\lambda\mod L$ are represented by $\lambda+L/pL$ so that the expression above equals
@@
Following the methodology of [SteinHecke], we derive explicit formulae for the action of Hecke operators on the Fourier expansion of vector-valued modular forms. In fact, the following is a modified version of [SteinHecke, Thm. 5.4 p. 246].
@@
+\;+ $p^{k-2}\delta_{\lambda\in{{}^{p}\!\mathcal{L}}}K_{L,p}g_{p}\left[1,{{}_{-}\chi_{p}}^{R},n-\operatorname{q}(p\cdot\ell_{\lambda/p})\right]\;a\left(\lambda,n\right)$
+\;+ $a(p\lambda,p^{2}n).$
Here, $K_{L,p}=p^{-m/2}G_{L,p}(1,1)$ and $g_{p}[r,\chi,n]$ are the Gauss sums from Definition 2.4, while $\ell_{\lambda/p}\in L^{\prime}$ is any element such that $\overline{p\ell_{p/\lambda}}=\lambda\in\mathcal{L}$. Further, ${{}_{p}\mathcal{L}}$ denotes the $p$-torsion of $\mathcal{L}$, ${{}^{p}\!\mathcal{L}}$ the $p$-multiples, and $\delta_{\lambda\in{{}^{p}\!\mathcal{L}}}=1$ if $\lambda\in{{}^{p}\!\mathcal{L}}$ and 00, otherwise. In addition, ${}_{-}\chi_{p}(h)=\left(\tfrac{-h}{p}\right)$ is the Legendre symbol and the exponent $R=R_{p}^{1}$ has been presented in Definition 2.6.
  • Proof:
@@
    Here, $\ell_{\mu}$ for some $\mu\in\mathcal{L}$ is understood to be an arbitrary but fixed lift of $\mu$ to $L^{\prime}$ and $\lambda/p$ is any fixed $p$-th root of $\lambda$ in $\mathcal{L}$. In ()($\ast$)( ∗ ), the dependency of the Gauss sum $G_{L,p}(1,-h)$ on $h$ may be extracted. In fact, an application of Remark 2.7 yields
@@
Proposition 4.17 suggests a potential approach for establishing bounds on the growth of Fourier coefficients of vector-valued modular forms by using the recursion relation for Fourier coefficients of eigenforms. For progressions in good primes, the procedure works straightforwardly, for bad primes, however, the situation is more delicate.
@@ 5.1. Definition and convergence
[…]
@@
[…]
  • Proof:
[…]
@@
For the purpose of deriving product expansions, we are interested in the case of one dimensional sublattices yielding the following definition.
@@
converges absolutely for $\Re(s)>k+1-2\sigma$ for some $21/4\leq\sigma\leq 1/$ 2 as follows.

### v2

@@ 3. Modular forms
Let $k\in(\mathbb{Z}/2)\setminus\mathbb{Z}$ with $k\geq 5/2$, $N\in\mathbb{N}$, and $\chi$ be a Dirichlet character modulo $4N$. Let
@@
be in $\mathcal{S}_{k}(\Gamma_{0}(4N),\chi)$. Then for $t\in\mathbb{N}$ and $\gcd(n,2N)=1$ we find
@@
The second bound goes back to Iwaniec [Iwa87] and Duke [Duk88] in a special case and has been generalised by Waibel [Wai18, Thm. 1 p. 186].
@@ 3.1. Eisenstein series
Beyond the theta series presented above, Eisenstein series represent another important class of modular forms. Our analysis requires the use of non-holomorphic Eisenstein series, which we introduce briefly based upon [Kie21, 2.5 pp. 30-36]. The statements are essential for our investigations, more specifically for deriving meromorphic continuation of the symmetric square type $L$-functions arising from the Rankin–Selberg type integrals in Section 5. In the following, let $(L,\operatorname{q})$ be a non-degenerate quadratic lattice and $k\in\tfrac{1}{2}\mathbb{Z}$ be a half integer and $\overline{\Gamma_{\infty}}\leq\operatorname{Mp}_{2}(\mathbb{Z})$ be the group generated by $\overline{T}$.
@@
The Eisenstein series $E_{L,\lambda,k}$ converges normally on $\mathbb{H}$ for $\Re(s)>1-\frac{k}{2}$, is real analytic in $\tau$ and an eigenfunction of the hyperbolic Laplace operator of weight $k$ with Eigenvalue s(s+k1)s(s+k-1).
@@
For $v\in\mathbb{R}^{\times}$ and $k\in\tfrac{1}{2}\mathbb{Z}$ representing the weight in the respective section, set
@@
where $\mathcal{W}_{\kappa,\mu}(z)$ is the usual Whittaker function as in [Olv+10, (13.14.3) p. 334].
Proposition 3.16 ([BK03, Prop. 3.1 p. 1695]).
@@
[…]
@@
According to [Hej83, p. 372] for an isotropic element $\lambda\in\mathcal{L}$
@@
From the Fourier expansion of $E_{L,\lambda,k}$ and the properties of the Whittaker function $\mathcal{W}_{s}$ (cf. [Olv+10, p. 335 (13.14.21)]) follows an asymptotic bound for the Eisenstein series which is relevant for the application of the Rankin-Selberg method.
@@
According to [Hej83, p. 372] we have the following transformation property.
@@ 4. Hecke theory
Note that this result is consistent with [Miy06, Lem. 4.5.7 (1) p. 140], where $T(p,p)=T_{p\mathcal{I}}$ and $T(1,p^{e})=\mathcal{T}(p^{e})$ in our notation. Note also that the number $N$ in Miyake may be assumed to equal $1$ in the current setting. By utilising the statement above, we may derive a relation for square indices which is the relevant case in the vector-valued setting.
@@
$=\;$ ${\mathcal{T}(p)}\left[\mathcal{T}(p)\mathcal{T}(p^{r})-p\cdot T_{p\mathcal{I}}\mathcal{T}(p^{r-1})\right]-p\cdot T_{p\mathcal{I}}\mathcal{T}(p^{r})$
$=\;$ $\left[\mathcal{T}(p^{2})+(p+1)\cdot T_{p\mathcal{I}}\right]\mathcal{T}(p^{r})$
@@
$=\;$ $\left[\mathcal{T}(p^{2})+(1-p)\cdot T_{p\mathcal{I}}\right]\mathcal{T}(p^{r})-(1+\delta_{r=2}p^{-1})\left(p\cdot T_{p\mathcal{I}}\right)^{2}\mathcal{T}(p^{r-2}).$
The Hecke operators introduced above not only constitute an abstract algebra but also operate on spaces of modular forms. Their action on Fourier coefficients of modular forms facilitates the derivation of product expansions for associated $L$-series, which motivates their examination in this work. The theory for scalar-valued modular forms is well known, however, naively extending the explicit construction of Hecke operators for the scalar-valued case to vector-valued modular forms fails, due to the Weil representation ${\rho}_{L}$ not being extendable to $\operatorname{GL}_{2}(\mathbb{Z}/\operatorname{lev}(L)\mathbb{Z})$ as a linear action. An alternative construction circumventing this issue is presented in [BS08] and is consistent with classical Hecke operators in the scalar-valued case. We will briefly sketch the content necessary for our proceedings.
@@
That the notation above is well-defined is a non-trivial fact (cf. [BS08, Sec. 5]). The same source established that the following operators define endomorphisms on $\mathcal{M}_{L,k}$.
@@
In the general case, by Definition 4.3 and the invariance of $f$ and $h$ under $\overline{\Gamma(1)}$, we may reduce to the choice $\alpha=g(m^{2})$. This, however is solved by the same argument as above combined with Lemma 4.5 in the components of the group algebra. ∎
@@
For any $M\in\mathbb{N}$, the algebra $\mathcal{H}_{M}^{2}$ acts on $\mathcal{S}_{L,k}$ as a commutative algebra of cusp form preserving, self-adjoint linear operators.
@@
Now that the existence of eigenforms is certain, a bound on the respective eigenvalues was advantageous. An effective method for obtaining such bounds was discovered by Kohnen [Koh87] even though the bound itself, at least for scalar-valued Siegel Modular forms, had been known before (cf. [Wei84, Sec. 6]).
@@
$=\;$ $\sum_{\lambda,\mu\in\mathcal{L}}\langle f_{\lambda}|_{k}g(m^{2})g^{\ast}(m^{2}),f_{\mu}|_{k}g(m^{2})g^{\ast}(m^{2})\rangle\langle\mathfrak{e}_{\lambda}|_{k}g(m^{2})g^{\ast}(m^{2}),\mathfrak{e}_{\mu}|_{k}g(m^{2})g^{\ast}(m^{2})\rangle$
$=\;$ $\sum_{\lambda,\mu\in\mathcal{L}}\langle f_{\lambda},f_{\mu}\rangle\lvert{{}_{m}\mathcal{L}}\rvert^{2}\langle\mathfrak{e}_{\lambda},\mathfrak{e}_{\mu}\rangle$
$=\;$ $\lvert{{}_{m}\mathcal{L}}\rvert^{2}\|f\|_{2}^{2}.$
@@
Finally, for the choice $m=p$, there are $p(p+1)$ cosets in the class $\overline{\Gamma(1)}\backslash\overline{\Gamma(1)}g(m^{2})\overline{\Gamma(1)}$ (cf. [Shi73, p. 451]), finishing the proof. ∎
@@ 4.1. Hecke action on Fourier coefficients
Note that for our purposes a formula that is valid for all primes is required. An advance in that direction has been made in [Ste15, Thm. 5.4 p. 246] for all odd primes. However, it has been pointed out by [BCJ18] that the cited Theorem is flawed. The original author has submitted a correction [Ste21], representing the previously arising, but flawed, coefficients in a more abstract fashion in terms of certain representation numbers of the lattice that also appear as coefficients of Eisenstein series in [BK01, p. 447].
These representation numbers, however, are not explicit enough to derive product expansions of $L$-series. Hence, the coefficients have to be calculated more explicitly based on the exposition [Ste15]. To this end we have to let $\mathcal{T}(p^{2})$ explicitly act on $f\in\mathcal{M}_{L,k}$ where the operation is given by letting right cosets of
@@
operate as in Definition 4.4. Recall that by [BS08, p. 263] or [Shi73, p. 451] we find
@@
The operation on the scalar component functions of the vector-valued modular form via the Petersson slash operator is well known and understood. As a consequence, it suffices to describe the action of the elements in (4.18) on the discriminant group. The action of $\gamma_{b}$ is presented in [Ste15, Thm. 5.1 p. 243] in greater generality than required.
@@
The action of $\beta_{h}$, however, will be extracted from [BCJ18, Prop. 5.3 p. 26] where we select the special case of $l=a=1$.
@@
where $G_{L,p}(n,h)=\sum_{v\in L/p^{n}L}e\left(\frac{h}{p^{n}}\operatorname{q}(v)\right)$ is the Gauss sum defined in Definition 2.4, $\ell_{\lambda/p}\in L^{\prime}$ is some element, such that $\cdot\overline{\ell_{\lambda/p}}=\lambda, p{{}^{p}\!\mathcal{L}}$ are the $p$-multiples in $\mathcal{L}$ and $\delta$ denotes a Kronecker symbol.
@@
where $\operatorname{q}(\mu)$ is to be understood as the quadratic form on $L^{\prime}$ of some preimage in $L^{\prime}$ of $\mu\in L^{\prime}/pL$. In order for the sum above not to vanish, it is necessary that $\lambda\in{{}^{p}\!\mathcal{L}}$ (cf. [Bar03, Lemma 5.2.1]). Note that by the isomorphism $(L^{\prime}/pL)/(L/pL)\simeq\mathcal{L}$, the solutions in $\mu$ of $\mu=\lambda\mod L$ are represented by $\lambda+L/pL$ so that the expression above equals
@@
Following the methodology of [Ste15], we derive explicit formulae for the action of Hecke operators on the Fourier expansion of vector-valued modular forms. In fact, the following is a modified version of [Ste15, Thm. 5.4 p. 246].
@@
$+\;$ $p^{k-2}\delta_{\lambda\in{{}^{p}\!\mathcal{L}}}K_{L,p}g_{p}\left[1,{{}_{-}\chi_{p}}^{R},n-\operatorname{q}(p\cdot\ell_{\lambda/p})\right]\;a\left(\lambda,n\right)$
$+\;$ $a(p\lambda,p^{2}n).$
Here, $K_{L,p}=p^{-m/2}G_{L,p}(1,1)$ and $g_{p}[r,\chi,n]$ are the Gauss sums from Definition 2.4, while $\ell_{\lambda/p}\in L^{\prime}$ is any element such that $\overline{p\ell_{p/\lambda}}=\lambda\in\mathcal{L}$. Further, $_{p}\mathcal{L}$} denotes the $p$-torsion of $\mathcal{L}$, $^{p}\!\mathcal{L}$} the $p$-multiples, and $\delta_{\lambda\in{{}^{p}\!\mathcal{L}}}=1$ if $\lambda\in{{}^{p}\!\mathcal{L}}$ and 0, otherwise. In addition, $_{-}\chi_{p}(h)=\left(\tfrac{-h}{p}\right)$ is the Legendre symbol and the exponent $R=R_{p}^{1}$ has been presented in Definition 2.6.
@@
Here, $\ell_{\mu}$ for some $\mu\in\mathcal{L}$ is understood to be an arbitrary but fixed lift of $\mu$ to $L^{\prime}$ and $\lambda/p$ is any fixed $p$-th root of $\lambda$ in $\mathcal{L}$. In $(\ast)$, the dependency of the Gauss sum $G_{L,p}(1,-h)$ on $h$ may be extracted. In fact, an application of Remark 2.7 yields
@@
Proposition 4.17 suggests a potential approach for estbalishing bounds on the growth of Fourier coefficients of vector-valued modular forms by using the recursion relation for Fourier coefficients of eigenforms. For progressions in good primes, the procedure works straightforwardly, for bad primes, however, the situation is more delicate.
@@ 5.1. Definition and convergence
For $f\in\mathcal{M}_{L,k}$, a splitting sublattice $M=L_{1}\oplus L_{2}\leq L$ with definite $L_{1}\otimes_{\mathbb{Z}}\mathbb{R}$, an isotropic element $\eta\in(L^{\prime}\cap L_{2}^{\prime})/(L\cap L_{2}^{\prime})$, and $t\in\mathbb{Q}^{\times}$ define the formal series
@@
Assume $L=L_{1}\oplus L_{2}$ splits with definite $L_{1}\otimes_{\mathbb{Z}}\mathbb{R}$ of rank $m_{1}$ and $2\leq k\in\tfrac{1}{2}\mathbb{Z}$. In case $f\in\mathcal{M}_{L,k}$, the series $L_{L_{1},\eta,r}(f,s)$ converges normally for $\Re(s)>k-1+\tfrac{m_{1}}{2}$, rendering it holomorphic. If, in addition, $f\in\mathcal{S}_{L,k}$, we find that $L_{L_{1},\eta,r}(f,s)$ converges normally for $\Re(s)>\frac{k+m_{1}}{2}-\sigma$ with $\sigma=1/2$ in the even case and $\sigma=1/4$ in the odd case.
By assumption $\sqrt{\lvert\operatorname{q}\rvert}:L_{1}\otimes_{\mathbb{Z}}\mathbb{R}\to\mathbb{R}_{\geq 0}$ defines a norm, rendering it equivalent to $\|\,\cdot\,\|_{\infty}$. This, together with the bound on Fourier coefficients from Corollary 3.13 reduces to bounding
@@
For the purpose of deriving product expansions, we are interested in the case of one dimensional sublattices yielding the following definition taylored to our setting.
@@
converges absolutely for $\Re(s)>k+1-2\sigma$ for some $1/4\leq\sigma\leq 1/2$ as follows.